OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4) I was thinkin 1/2 but i dont kno

OpenStudy (anonymous):

\[\frac{ 4-2 }{ 7-6} = ?\] Can you go from there?

OpenStudy (anonymous):

nope, i dont kno how to do this

OpenStudy (anonymous):

tbh i jus need the answer

OpenStudy (anonymous):

You subtract 2 from 4, and you get 2. Then, you subtract 6 from 7, and get 1. Therefore, your slope is 2/1. Do you get it? Remember, this is the formula to find the slope: \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

yeah i was gonna put that but i was scared i was wrong, lol

OpenStudy (anonymous):

so would the answer be -2

OpenStudy (anonymous):

No, it is 2/1, or 2.

OpenStudy (anonymous):

oh ok, im sorry that slipped my mind but thank you so much

OpenStudy (anonymous):

No problem, glad I was of help

OpenStudy (anonymous):

you helped very much